countryboy:

A hose dispenses 1,413 cm³ of water every minute into a cone with a radius of 60 cm and a height of 150 cm. How long will it take to fill the cone? Use 3.14 to approximate pi. Enter your answer in the box. min

jayrules232:

Volume of the cone: V = 1/3 r^2 Pi * h. V = 1/3 * 60^2 * 3.14 * 150 ; V = 1/3 * 3600 * 471; V = 565,200 cm^3. If 1413 cm^3 of water goes every minute into the cone, we have to divide the volume by 1413: 565,200 : 1413 = 400. Answer: It will take 400 minutes.

Shadow:

@countryboy Do you need further help?